shift summary for evening shift 8/July/99

Jim Hirschauer (jhirsch@cuhep.Colorado.EDU)
Thu, 8 Jul 1999 23:07:51 -0600 (MDT)

Shift Summary for evening shift, 8/July/99

On Shift: Viv O'Dell, Jim Hirschauer

Hours of beam delivered at all intensities: 8.5
Hours of beam delivered at nominal intensity: 8.5
Hours of E832 data (special runs not included): 8.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

========================================================================
Runs ended this shift: Run 13838 - in progress - BEAM_832

Tapes brought to FCC this shift: QKC098,QKC099,QKC100
QKS098,QKS099,QKS100
QKK098,QKK099,QKK100
QKB038

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Smooth E832 datataking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13838 15:40 Confusion over SEM calibration. TV screen doesn't
agree w/ the scaler.
TV, epicure - 4.16E12
scaler - 3.45E12

15:50 R. Ford says Sdaq NM2 is correct, and CH 13 is
wrong. Top people from MCR are working on it.
They will also check to see if all SEM's add up,
Rick says his calibrations are a week old and may
have drifted slightly.

16:30 We noticed the HCC inefficiency is very poor in
some channels. Elliott looked at the frowny faces
for awhile and said he had top people on it.

16:50 RF called back. CH13 NM2 SEM now agrees w/ slow
daq, but there is an overall matter
non-conservation effect. It may be a) our calib
b) TeVatron calib c)matter is being created
Top people will investigate tomorrow

260 18:15 Hogan adjusted variable filter wheel by 2000 ticks
during spill 260. Laser light level -> 2X

23:15 Notice hot bit in HCC plot - Channel #2622

00:00 Hand over to Julie and Miguel, run 13838 in
progress and running smoothly

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

o keep collecting BEAM_832 data - no Muon run